Let two cars have the same mass m. Suppose a car passes through a stop sign with out stopping at a speed of 5 mi/hr. Another car travels from the west at at speed of 70 mi/hr. The two cars stick together and travel off the road at some final speed v and the angle θ as measured relative to the +x−axis.
Answer #1

Assuming first car ti travelling towards +y axis or towards north

conservation of momentum

in x direction

m*70=2*m*Vx

Vx=35 mi/hr

in y direction

m*5=2*m*Vy

Vy=2.5 mi/hr

θ=arctan(Vy/Vx)

=4.0856